# Hybrid classical-quantum communication networks

### v1

@@ 3 Integration of quantum communications in conventional fiber-optic networks
Early quantum communication experiments were largely conducted over “dark fiber”, where no classical communication signals were present alongside quantum signals. The rationale is clear: classical signals are many orders of magnitude more intense than quantum signals, and even minor leakage from classical signals can severely degrade quantum signals. However, dark-fiber links are scarce and expensive. Given the limited near-term applications of quantum networking, enabling the coexistence of quantum and classical traffic within standard WDM fiber networks offers a promising route to commercialization. To this end, significant research has focused on enhancing the compatibility of quantum communication protocols with existing fiber-optic infrastructures. A detailed review of the coexistence of QKD and classical traffic in the same fiber can be found in [55].
@@ 3.1 Coexistence of quantum and classical signals in telecom fiber
Indeed, the initial drivers for WDM in quantum communications centered on the coexistence of QKD signals with classical traffic, with the first demonstration placing the former in the O-band (1260–1360 nm) and the latter in the C-band (1530–1565 nm) [232]. A quantum/classical O/C-band split has been the preferred allocation in most subsequent QKD coexistence experiments [51, 53, 255, 162, 243, 87, 32], although other CWDM channel combinations—such as quantum in the C-band and classical in L-band (1565–1625 nm) [174], or quantum in the C-band and classical in the O-band, L-band, and S-band (1460–1530 nm) [78]—have been realized as well. Pushing even tighter to DWDM spacings has enabled positive QKD rates over tens of km with copropagating classical channels as close as 200 GHz [184, 76], highlighting the feasibility of ultradense bandwidth allocations.
More recently, quantum-classical coexistence experiments have branched out beyond QKD to more general quantum characterization scenarios leveraging quantum process  [42] and quantum state [230] tomography. Arguably the defining impairment in all of these experiments is Raman scattering—noise photons generated through light interactions with vibrational modes in silica fiber [217]. Related to both the classical launch power and fiber length, Raman is challenging to handle, since its spectral content can overlap with the quantum band of interest; the fact its intensity is slightly higher on the redshifted side of the signal and drops for large spectral separations explains the general experimental preference for O-band quantum signals paired with C-band classical ones, yet the DWDM experiments above [184, 76, 42]—as well as nonlocal biphoton modulation with coexisting radio-over-fiber timing synchronization signals [50]—highlight the feasibility for C-band coexistence as well, where both signals can enjoy the lowest-loss region of optical fiber.
@@ 3.2 Quantum incompatible devices in classical network and solutions
Classical networks are carefully engineered such that the transmitted data signals are reliably received by the paired receivers. This requires that there be sufficient SNR at the receiver to faithfully decode the transmitted classical information. In addition, with current technology, it is common that classical receivers are insensitive to quantum signals. Consider a quantum transmitter which emits a a stream of perfect 1550-nm single photons (each about 0.8 eV) at 100 billion times a second. Such a quantum transmitter, while beyond the realm of current experimental feasibility, would have an average launch power of roughly four orders of magnitude below that of a typically launched conventional signal (roughly 1 mW or 0 dbm). In addition, classical network devices can emit optical noise that is below the sensitivity threshold of classical receivers. As a result, conventional network devices may not emit noise detectable by a classical receiver, yet they may emit significant noise for quantum networks, potentially impairing the function of a quantum receiver.
One such device is an optical amplifier. While in the context of no-cloning, an amplifier adds noise to amplified photons, it can also add noise to a quantum network without an input signal via amplified spontaneous emission. ASE can have a broad bandwidth, and care must be take to prevent that noise from being introduced directly into a quantum network. For example, the most common types of networked amplifiers, EDFA, have ASE bandwidths from 1530 nm - 1565 nm. Other similar amplifiers include Silicon Optical Amplifiers (SOAs), which also emit ASE noise. Semiconductor lasers can emit noise at much higher bandwidths than the emission wavelength albeit at a much lower intensity than the fundamental emission wavelength. If the spontaneous emission overlaps with the quantum signal wavelength(s), that part of the spontaneous nose spectrum may need to be filtered out before the spectrum is combined onto a fiber with a quantum signal. In recognition of these noise sources, an early experiment used an amplifier for mid-span amplification of conventional signals in the C-band, while bypassing an O-band quantum signal[167]. In addition, both the quantum signal and conventional laser light were filtered to reduce noise in early QKD coexistence experiments[184].
@@ 3.3 Simultaneous transmission of classical and quantum information
A significant next step in this direction is the shared use of network resources, including both hardware and software, for classical and quantum communications. In particular, CV QKD is especially attractive because its information encoding and decoding methods, as well as the required hardware, are remarkably similar to those of classical coherent optical communication. Recent advancements in both fields have further reduced the gap between them. On the classical side, continuous improvements in detector performance and the development of forward error correction (FEC) have enabled bit error rates (BER) as low as $10^{-9}$ with just a few photons per bit [207]. This implies that state-of-the-art coherent communication systems can already operate in the quantum noise-dominant regime. On the quantum side, new security proof technologies in CV QKD [83, 144] allow much-simplified encoding schemes, such as binary phase-shift keying (BPSK) and quadrature phase-shift keying (QPSK), which are widely used in classical communication. Furthermore, novel phase recovery methods [190, 205], now enable CV QKD to use a true LO generated at the receiver, eliminating the need for a transmitted LO from the sender.
@@ 3.4 Beyond the physical layer
Yet despite the potential value of a quantum Internet stack, nearly all quantum communications research so far has been confined to what the TCP/IP stack would place in the physical layer, namely, protocols for and transmission of quantum information over point-to-point links. Of course, the unique features of quantum communications derive from the physical properties of quantum systems, so the historical emphasis on the physical layer was not only reasonable, but required to push the field forward. However, with growing interest in more general multinode quantum networks, several quantum stacks analogous to TCP/IP have been proposed. The quantum recursive network architecture represents one of the earliest formulations with explicitly quantum layers [160]. Other proposed stacks focus on the establishment of multipartite graph states [173], distinguish layers by the increasing reliability and distance of entanglement generation [127, 65, 113], or identify layers with functionalities in wavelength-multiplexed entanglement distribution [19, 17]. The latter two stacks have even been implemented experimentally in deployed testbeds with nitrogen-vacancy [175] or photonic polarization [19] qubits, respectively.
@@
The ways in which packet and circuit switching impact SDN highlight the generally important relationship between SDN and the adopted quantum Internet stack. For just as classical SDN provides resources for managing a data plane in the TCP/IP stack, the types of routing decisions made by a quantum-conventional SDN controller will depend on the architecture of the quantum data plane; the stacks proposed above offer provisional examples for continued research [160, 173, 127, 65, 113, 19, 17]. Therefore while the natural synergy of SDN with quantum networking sparks optimism in the path toward integrating quantum and classical networks on large scales, much work remains–both in fleshing out detailed protocols and in developing the technology capable of realizing them.
@@ 3.5 Quantum network demonstrations
The ultimate goal of any network—quantum or classical—is to enable communications between distributed users, and thus are often classified by the spatial scale of the nodes serviced: e.g., local, metro, and wide area networks (LANs, MANs, and WANs, respectively) and their quantum equivalents (QLANs, QMANs, and QWANs). Apart from the specific application of connecting modular quantum computes in a single laboratory, quantum networks by their very nature must move from the lab to the field to realize their potential. In this section we overview quantum network testbeds that have made this important jump. In line with the overall scope of this review, we focus on fiber-optic testbeds rather than free-space, yet we note the importance of satellite quantum communications for long-distance networking, particularly in the current pre-repeater era. Of multiple efforts globally [214], the Micius mission has realized especially groundbreaking experimental demonstrations of satellite-to-ground QKD and entanglement distribution; we recommend a recent review [132] for a summary of the progress so far.
When transferring a fiber-optic quantum communications experiment from the tabletop to a deployed network, the key challenges can be classified in two broad categories: (i) those related to the quantum channel, and (ii) those related to node synchronization. Examples of (i) include temperature variations, strain, and even air currents subjected to deployed fiber that a fiber spool in a well-controlled lab does not face, while time-stamping detection events and synchronizing coherent operations between distributed sites represent examples of (ii). A common strategy in quantum communication experiments considers loopbacks in which the send and receive connections of a deployed fiber are terminated in the same location; while this arrangement is sensitive to real-world channel impairments, it sidesteps the synchronization problem altogether, giving all ostensive nodes access to the same local clock for modulation and coincidence counting [^1] Such loopbacks are distinct from bidirectional use of a single fiber, as in plug-and-play QKD, where Alice attenuates and reflect light back to Bob. In this case, Alice must synchronize her modulation modulation signal via photodetection [210].. Given this significant simplification for loopbacks, we concentrate our review on testbeds with truly distributed sites, as validated by the measurement of entanglement or generation of a secret key in separate labs. For the sake of consistency, we report all distances from prior experiments in terms of deployed fiber lengths, excluding any extensions from additional spools. Note that many experiments referenced below also mention geographic distance—a number of particular importance in closing the locality loophole in Bell tests—which can differ significantly from the in-fiber number.
@@
Despite the diversity in geography and applications of each of these testbeds, several trends have emerged. First, although entanglement-based QKD [77, 181, 209, 69] and CV-QKD [181, 139] have been considered in several of them, decoy-state BB84 with phase or polarization encoding has proven by far the most common protocol implemented, a testament to both its simplicity and performance. Second, nearly all testbeds have leveraged some form of wavelength-multiplexed classical pulses (CWDM or DWDM) for clock recovery and timing synchronization. And finally, all have supported fiber link distances on the order of $\sim$10–100 km, with trusted nodes enlisted to extend to distances beyond those of elementary links. Indeed, pushing the trusted node paradigm even further has spawned true QWANs that connect multiple QMANs, first joining the Hefei and Wuhu QKD networks [244] and more recently uniting QMANs in Beijing, Jinan, Hefei, and Shanghai [54]. In the latter case, nodes as far as 2000 km apart in fiber (4600 km including the satellite extension) can establish secret keys—far and away the longest distance of any deployed network.
@@
Entanglement distribution in deployed fiber was demonstrated first in Geneva, with time-bin-entangled photons distributed from a single source to two locations through 8.1 km and 9.3 km fiber channels [224]. In both this and a subsequent nonlocality test [204], coincidences were obtained via classical transmission of the detector output signals back to the source, thus supporting spatially distributed detection without explicit time synchronization. Other early work distributing polarization entanglement in deployed fiber appeared in Vienna [177, 233]; enlisted for entanglement-based QKD, this system complemented the coherent-state systems also supported in the larger QKD testbed [181]. Recent work has pushed polarization entanglement distribution to fiber separations of 96 km in submarine [250] and 248 km in terrestrial [166] links, while time-bin-entanglement has recently been measured over 50 km in deployed fiber [183].
While most testbeds are focused on the distribution of DV entanglement, CV networks could prove valuable, especially in the context of quantum sensor networks. In particular, distributed multi-mode squeezing could be a useful network resource. To this end, part of the Oak Ridge quantum network distributed two-mode squeezing of over 1.5 km of deployed optical fiber in the C-band[45].
While photonic quantum teleportation can be enabled with DV or CV resources, it has not yet attained the same distances as elementary entanglement distribution. Time-bin qubits have nonetheless been teleported 0.8 km [145], 17.3 km [242], and 30.4 km [218] in terms of total deployed fiber distance from Alice to Bob. On the matter qubit front, nitrogen-vacancy (NV) centers separated by 1.7 km of deployed fiber were entangled in one of the seminal loophole-free Bell tests [91], with a recent extension to 25 km enabled by quantum frequency conversion [222]; atomic ensembles have been entangled over 17 km of deployed fiber [150], as well as individual atoms separated by 700 m [266, 240]; ion-ion [115] and ion-photon [117] entanglement has been realized over 520 m and 14.4 km, respectively; a photon entangled with a rare-earth quantum memory has been measured after 44 km [197]; and two SiV spin qubits connected by a 35 km loopback (though located in separate rooms) have been entangled as well [125].
@@
Testbeds throughout the world are pushing toward the multinode regime, and two in particular—at Delft University of Technology [178, 175] and Oak Ridge National Laboratory (ORNL) [19, 17, 9, 1]—have recently confirmed distributed entanglement. The Delft testbed is based on NV center qubits with a fiber span of 30 m; the Oak Ridge testbed is based on photonic qubits only, but is directly compatible with the lightwave communications infrastructure: operating in the C-band, using WSSs, and including a 1.2 km link separation in standard SMF. Both rely on quantum stacks analogous to TCP/IP and have explored the interplay between state quality and state quantity, via either a fidelity/latency tradeoff [178, 175] or distillable entanglement [19, 17, 9, 1]. Such network metrics highlight the desire to move beyond QKD to more general network applications in sensing and computing; indeed, demonstrations of remote state preparation on both networks [19, 175] point specifically to applications in blind quantum computing [28, 30, 79].
@@
There remain many directions for improvement on existing networks like that in Fig. 8. In addition to expanding distance after the advent of quantum repeaters, an important target is the integration of matter qubits with a telecom fiber infrastructure (essentially combining the strengths of the Delft and Oak Ridge testbeds), either through quantum frequency converters and transducers [126, 143, 3, 95] or fully telecom-compatible photonic storage [215, 128]. From an application perspective, a multinode network demonstration of quantum-enhanced capabilities other than QKD—such as entanglement-enhanced sensing, distributed quantum computing, or blind quantum quantum cloud computing—would mark a critical milestone in the field, potentially heralding an era of quantum networks that are valued primarily for their practical value than their research interest.

### v2

@@ 3 Integration of quantum communications in conventional fiber-optic networks
Early quantum communication experiments were largely conducted over “dark fiber”, where no classical communication signals were present alongside quantum signals. The rationale is clear: classical signals are many orders of magnitude more intense than quantum signals, and even minor leakage from classical signals can severely degrade quantum signals. However, dark-fiber links are scarce and expensive. Given the limited near-term applications of quantum networking, enabling the coexistence of quantum and classical traffic within standard WDM fiber networks offers a promising route to commercialization. To this end, significant research has focused on enhancing the compatibility of quantum communication protocols with existing fiber-optic infrastructures. Table 3 summarizes the major challenges associated with transmitting quantum signals over classical fiber-optic networks, along with potential mitigation strategies. A detailed review of the coexistence of QKD and classical traffic in the same fiber can be found in [115].
Table 3: Transmitting quantum signals through classical fiber-optic networks.
Challenges Possible mitigation strategies
Noise photons generated by intense classical traffic can significantly degrade the SNR of quantum communications (including crosstalk and scattered conventional signal noise). (a) Optimize wavelength assignment in WDM networks to minimize crosstalk [197]. (b) Apply additional spectral and temporal filtering—such as narrowband optical filters and detector gating—to suppress noise photons in quantum channel [161]. (c) Leverage the intrinsic mode-selectivity of optical homodyne detectors in coherent quantum communication to reject out-of-mode noise [20]. (d) Use novel fiber types, such as hollow-core fibers, which exhibit reduced nonlinearity and scattering effects, thereby minimizing classical-to-quantum channel interference [198]. (e) Use Procrustean local filtering to remove conventional signal crosstalk from distributed entanglement [199].
Certain devices commonly used in classical optical networks—such as optical amplifiers, classical repeaters, and routers—are incompatible with quantum communications. (a) Bypass devices that are incompatible with quantum signals  [200]. Note that such bypasses must block amplified spontaneous emission or other noise emitted by the bypassed device, including amplifiers and lasers [201], in the quantum transmission band. (b) Develop and deploy hybrid network components, such as quantum-compatible routers, that can support both classical and quantum communications [150].
Quantum communications are generally more sensitive to channel impairments and noise than classical systems. Advance noise-reduction technologies—such as precise polarization control, dispersion management, synchronization, and optical phase recovery—which enhance the performance of quantum communication and can also benefit classical channels.
@@ 3.1 Coexistence of quantum and classical signals in telecom fiber
Historically, the initial drivers for WDM in quantum communications centered on the coexistence of DV-QKD signals with classical traffic, with the first demonstration placing the former in the O-band (1260–1360 nm) and the latter in the C-band (1530–1565 nm) [197]. A quantum/classical O/C-band split has been the preferred allocation in most subsequent QKD coexistence experiments [162, 175, 202, 203, 204, 205, 206], although other CWDM channel combinations—such as quantum in the C-band and classical in L-band (1565–1625 nm) [207], or quantum in the C-band and classical in the O-band, L-band, and S-band (1460–1530 nm) [208]—have been realized as well. Pushing even tighter to DWDM spacings has enabled positive QKD rates over tens of kilometers with copropagating classical channels as close as 200 GHz [161, 209], highlighting the feasibility of ultradense bandwidth allocations.
More recently, quantum-classical coexistence experiments have branched out beyond QKD to more general scenarios such as quantum process tomography [210], quantum state tomography [211], and quantum teleportation [212]. Arguably the defining impairment in all of these experiments is Raman scattering—noise photons generated through light interactions with vibrational modes in silica fiber [213]. Related to both the classical launch power and fiber length, Raman is challenging to handle, since its spectral content can overlap with the quantum band of interest; the fact its intensity is slightly higher on the redshifted side of the signal and drops for large spectral separations explains the general experimental preference for O-band quantum signals paired with C-band classical ones, yet the DWDM experiments above [161, 209, 210]—as well as nonlocal biphoton modulation with coexisting radio-over-fiber timing synchronization signals [214]—highlight the feasibility for C-band coexistence as well, where both signals can enjoy the lowest-loss region of optical fiber.
@@ 3.2 Quantum incompatible devices in classical network and solutions
Classical networks are carefully engineered such that the transmitted data signals are reliably received by the paired receivers. This requires that there be sufficient SNR at the receiver to faithfully decode the transmitted classical information. In addition, with current technology, it is common for classical receivers to be completely insensitive to quantum signals. Consider a quantum transmitter which emits a stream of perfect 1550 nm single photons (each about 0.8 eV) at 100 billion times a second. Such a quantum transmitter, while beyond the realm of current experimental feasibility, would have an average launch power of roughly five orders of magnitude below that of a typically launched conventional signal (roughly 1 mW or 0 dBm). In addition, classical network devices can emit optical noise that is below the sensitivity threshold of classical receivers. As a result, conventional network devices may not emit noise detectable by a classical receiver, yet they may emit significant noise for quantum networks, potentially impairing the function of a quantum receiver.
One such device is an optical amplifier. While in the context of no-cloning, an amplifier adds noise to amplified photons, it can also add noise to a quantum network without an input signal via ASE. ASE can have a broad bandwidth, and care must be take to prevent that noise from being introduced directly into a quantum network. For example, the most common types of networked amplifiers, EDFA, have ASE bandwidths from 1530–1565 nm. Other similar amplifiers include semiconductor optical amplifiers (SOAs), which also emit ASE noise. Semiconductor lasers can emit noise at much higher bandwidths than the emission wavelength albeit at a much lower intensity than the fundamental emission wavelength. If the spontaneous emission overlaps with the quantum signal wavelength(s), that part of the spontaneous nose spectrum may need to be filtered out before the spectrum is combined onto a fiber with a quantum signal. In recognition of these noise sources, an early experiment used an amplifier for mid-span amplification of conventional signals in the C-band, while bypassing an O-band quantum signal [200]. In addition, both the quantum signal and conventional laser light were filtered to reduce noise in early QKD coexistence experiments [161].
@@ 3.3 Simultaneous transmission of classical and quantum information
A significant next step in this direction is the shared use of network resources, including both hardware and software, for classical and quantum communications. In particular, CV-QKD is especially attractive because its information encoding and decoding methods, as well as the required hardware, are remarkably similar to those of classical coherent optical communication. Recent advancements in both fields have further reduced the gap between them. On the classical side, continuous improvements in detector performance and the development of forward error correction (FEC) have enabled bit error rates (BER) as low as $10^{-9}$ with just a few photons per bit [215]. This implies that state-of-the-art coherent communication systems can already operate in the quantum noise-dominant regime. On the quantum side, new security proof technologies in CV-QKD [216, 217] allow much-simplified encoding schemes, such as binary phase-shift keying (BPSK) and quadrature phase-shift keying (QPSK), which are widely used in classical communication. Furthermore, novel phase recovery methods [125, 126], now enable CV-QKD to use a true LO generated at the receiver, eliminating the need for a transmitted LO from the sender. This local LO scheme has been experimentally demonstrated over 100 km of optical fiber [218].
@@ 3.4 Beyond the physical layer
Yet despite the potential value of a quantum Internet stack, nearly all quantum communications research so far has been confined to what the TCP/IP stack would place in the physical layer, namely, protocols for and transmission of quantum information over point-to-point links. Of course, the unique features of quantum communications derive from the physical properties of quantum systems, so the historical emphasis on the physical layer has not only been reasonable, but required to push the field forward. However, with growing interest in more general multinode quantum networks, several quantum stacks analogous to TCP/IP have been proposed. The quantum recursive network architecture represents one of the earliest formulations with explicitly quantum layers [226]. Other proposed stacks focus on the establishment of multipartite graph states [227], distinguish layers by the increasing reliability and distance of entanglement generation [228, 229, 230], or identify layers with functionalities in wavelength-multiplexed entanglement distribution [158, 231]. The latter two stacks have even been implemented experimentally in deployed testbeds with nitrogen-vacancy (NV) [232] or photonic polarization [158] qubits, respectively.
@@
The ways in which packet and circuit switching impact SDN highlight the generally important relationship between SDN and the adopted quantum Internet stack. For just as classical SDN provides resources for managing a data plane in the TCP/IP stack, the types of routing decisions made by a quantum-conventional SDN controller will depend on the architecture of the quantum data plane; the stacks proposed above offer provisional examples for continued research [226, 227, 228, 229, 230, 158, 231]. Therefore while the natural synergy of SDN with quantum networking sparks optimism in the path toward integrating quantum and classical networks on large scales, much work remains—both in fleshing out detailed protocols and in developing the technology capable of realizing them.
@@ 3.5 Quantum network demonstrations
The ultimate goal of any network—quantum or classical—is to enable communications between distributed users, and thus networks are often classified by the spatial scale of the nodes serviced: e.g., local, metro, and wide area networks (LANs, MANs, and WANs, respectively) and their quantum equivalents (QLANs, QMANs, and QWANs). Apart from the specific application of connecting modular quantum computers in a single laboratory, quantum networks by their very nature must move from the laboratory to the field to realize their potential. In this section we overview quantum network testbeds that have made this important jump. In line with the overall scope of this review, we focus on fiber-optic testbeds rather than free-space, yet we note the importance of satellite quantum communications for long-distance networking, particularly in the current pre-repeater era. Of multiple efforts globally [240], the Micius mission has realized especially groundbreaking experimental demonstrations of satellite-to-ground QKD and entanglement distribution; we recommend a recent review [241] for a summary of the progress so far.
When transferring a fiber-optic quantum communications experiment from the tabletop to a deployed network, the key challenges can be classified in two broad categories: (i) those related to the quantum channel, and (ii) those related to node synchronization. Examples of (i) include temperature variations, strain, and even air currents subjected to deployed fiber that a fiber spool in a well-controlled laboratory does not face, while time-stamping detection events and synchronizing coherent operations between distributed sites represent examples of (ii). A common strategy in quantum communication experiments considers loopbacks in which the send and receive connections of a deployed fiber are terminated in the same location; while this arrangement is sensitive to real-world channel impairments, it sidesteps the synchronization problem altogether, giving all ostensive nodes access to the same local clock for modulation and coincidence counting [^1] Such loopbacks are distinct from bidirectional use of a single fiber, as in plug-and-play QKD, where Alice attenuates and reflects light back to Bob. In this case, Alice must synchronize her modulation modulation signal via photodetection [242].. Given this significant simplification for loopbacks, we concentrate our review on testbeds with truly distributed sites, as validated by the measurement of entanglement or generation of a secret key in separate labs. For the sake of consistency, we report all distances from prior experiments in terms of deployed fiber lengths, excluding any extensions from additional spools. Note that many experiments referenced below also mention geographic distance—a number of particular importance in closing the locality loophole in Bell tests—which can differ significantly from the in-fiber number.
@@
Despite the diversity in geography and applications of each of these testbeds, several trends have emerged. First, although entanglement-based QKD [244, 106, 108, 110] and CV-QKD [106, 109] have been considered in several of them, decoy-state BB84 with phase or polarization encoding has proven by far the most common protocol implemented, a testament to both its simplicity and performance. Second, nearly all testbeds have leveraged some form of wavelength-multiplexed classical pulses (CWDM or DWDM) for clock recovery and timing synchronization. And finally, all have supported fiber link distances on the order of $\sim$10–100 km, with trusted nodes enlisted to extend to distances beyond those of elementary links. Pushing the trusted node paradigm even further has spawned true QWANs that connect multiple QMANs, first joining the Hefei and Wuhu QKD networks [249] and more recently uniting QMANs in Beijing, Jinan, Hefei, and Shanghai [250]. In the latter case, nodes as far as 2000 km apart in fiber (4600 km including the satellite extension) can establish secret keys—far and away the longest distance of any deployed network.
@@
Entanglement distribution in deployed fiber was demonstrated first in Geneva, with time-bin-entangled photons distributed from a single source to two locations through 8.1 km and 9.3 km fiber channels [251]. In both this and a subsequent nonlocality test [252], coincidences were obtained via classical transmission of the detector output signals back to the source, thus supporting spatially distributed detection without explicit time synchronization. Other early work distributing polarization entanglement in deployed fiber appeared in Vienna [253, 173]; enlisted for entanglement-based QKD, this system complemented the coherent-state systems also supported in the larger QKD testbed [106]. Recent work has pushed polarization entanglement distribution to fiber separations of 96 km in submarine [254] and 248 km in terrestrial [255] links, while time-bin entanglement has recently been measured over 50 km in deployed fiber [256].
While most testbeds are focused on the distribution of DV entanglement, CV networks could prove valuable, especially in the context of quantum sensor networks. In particular, distributed multi-mode squeezing could be a useful network resource. To this end, part of the Oak Ridge quantum network (discussed below) has distributed two-mode squeezing of over 1.5 km of deployed optical fiber in the C-band [257].
While photonic quantum teleportation can be enabled with DV or CV resources, it has not yet attained the same distances as elementary entanglement distribution. Time-bin qubits have nonetheless been teleported 0.8 km [258], 17.3 km [259], and 30.4 km [260] in terms of total deployed fiber distance from Alice to Bob. On the matter qubit front, NV centers separated by 1.7 km of deployed fiber were entangled in one of the seminal loophole-free Bell tests [90], with a recent extension to 25 km enabled by quantum frequency conversion [261]; atomic ensembles have been entangled over 17 km of deployed fiber [262], as well as individual atoms separated by 700 m [94, 92]; ion-ion [263] and ion-photon [264] entanglement has been realized over 520 m and 14.4 km, respectively; a photon entangled with a rare-earth quantum memory has been measured after 44 km [265]; and two SiV spin qubits connected by a 35 km loopback (though located in separate rooms) have been entangled as well [93].
@@
Testbeds throughout the world are pushing toward the multinode regime. For the purposes of this review, we focus on two in particular—at Delft University of Technology [266, 232, 261] and Oak Ridge National Laboratory (ORNL) [158, 231, 267, 159, 268]—that have pioneered distributed multinode entanglement [^2] As three-node entanglement-based demonstrations are appearing in the literature at a rapid pace [269, 270, 271], our intent with the Delft-ORNL focus is not to exclude other leading testbeds, but rather highlight important engineering considerations relevant to hybrid classical-quantum networks as a whole.. The Delft testbed is based on NV center qubits; the Oak Ridge testbed is based on photonic qubits only, but is directly compatible with the lightwave communications infrastructure: operating in the C-band, using WSSs, and including a 1.2 km link separation in standard SMF. Both rely on quantum stacks analogous to TCP/IP and have explored the interplay between state quality and state quantity, via either a fidelity/latency tradeoff [266, 232] or distillable entanglement [158, 231, 267, 159]. Such network metrics highlight the desire to move beyond QKD to more general network applications in sensing and computing; indeed, demonstrations of remote state preparation (RSP) on both networks [158, 232] point specifically to applications in blind quantum computing [272, 273, 274].
@@
There remain many directions for improvement on existing networks like that in Fig. 8. In addition to expanding distance after the advent of quantum repeaters, an important target is the integration of matter qubits with a telecom fiber infrastructure. Central to such a vision are quantum transducers, devices that coherently convert quantum information from one physical system to another [179, 180, 181]. Ultimately, it is desirable that such transducers would function fault tolerantly. We are excited by the sustained improvement in quantum transducers spanning the particularly difficult microwave-to-optical energy regimes, leveraging physical mechanisms ranging from optomechanical [286] and electro-optic resonators [287] to neutral atoms [288] and ions [289]. To the extent quantum memories supporting direct telecom-band photon storage are available—also an active research front [290, 291]—transduction between widely different energies could potentially be averted, at least in the core of the optical network.
From an application perspective, a multinode network demonstration of quantum-enhanced capabilities other than QKD—such as entanglement-enhanced sensing [292, 293, 294, 295], entanglement-enhanced telescopy [296, 297], distributed quantum computing [298, 299], or blind quantum quantum cloud computing [272]—would mark a critical milestone in the field, potentially heralding an era of quantum networks that are valued primarily for their practical value than their research interest.
Table 4 concludes this subsection with a summary of key entanglement distribution experiments mentioned above. In keeping with this review’s deployed focus, we include only experiments in which the entangled particles are measured nonlocally (i.e., without looping back into the same laboratory) and define the maximum distance as the length of the fiber lightpath between the farthest entangled nodes. In our view, this distance definition is most relevant to hybrid classical-quantum networking, where we typically envision building a quantum network on top of existing optical fiber infrastructure (nodes and fiber paths). In such a context, the shortest “straight-line” distance is not generally available, and additional fiber spools—often inserted to add delay or simulate longer links—unnecessarily reduce performance in a application-focused network. Entanglement is classified by particle type: flying photonic qubits or stationary memory qubits (e.g., color centers, atoms, or ions). The most advanced protocol implemented in each reference comprises the fifth column, from baseline entanglement distribution to entanglement swapping or an application like QKD.
Analogous to the Internet Protocol itself, we have curated Table 4 as a “best effort” service; no guarantee of exhaustivity is made. Nevertheless, perhaps the most striking feature is the explosion of results in the last three years. Although covering a 26-year span (1998–2024), more than half of all references in Table 4 have appeared since 2022, providing yet another indication of the growing interest in quantum networking globally.
Table 4: Summary of entanglement distribution experiments in deployed fiber.
# Max. fiber Year
Location nodes distance [km] Entanglement Experiment published
Geneva 3 17.4 photon/photon Bell test 1998 [251]
Vienna 2 1.45 photon/photon QKD 2004 [253]
Geneva 2 0.8 photon/photon teleportation 2007 [258]
Geneva 3 30.9 photon/photon Bell test 2008 [252]
Vienna 2 16 photon/photon QKD 2009 [173]
Delft 3 1.7 memory/memory Bell test 2015 [90]
Hefei 3 30.4 photon/photon teleportation 2016 [260]
Calgary 3 17.3 photon/photon teleportation 2016 [259]
Hefei 3 25.8 photon/photon swapping 2017 [300]
Malta–Sicily 2 96 photon/photon distribution 2019 [254]
Delft 3 0.03 memory/memory RSP 2021 [266]
Oak Ridge 3 1.5 photon/photon RSP 2021 [158]
Munich 3 0.7 memory/memory swapping 2022 [92]
Munich 2 0.7 memory/memory QKD 2022 [94]
Hefei 2 17 memory/memory distribution 2022 [262]
Austria–Slovakia 3 248 photon/photon distribution 2022 [255]
Innsbruck 2 0.52 memory/memory swapping 2023 [263]
Nice 3 50 photon/photon QKD 2023 [256]
Barcelona 2 44 photon/memory distribution 2023 [265]
Oak Ridge 3 1.5 photon/photon digital signatures 2024 [301]
DC Metro 3 0.25 photon/photon distribution 2024 [269]
Hefei 3 21.6 memory/memory swapping 2024 [270]
Oak Ridge 6 2.1 photon/photon distribution 2024 [268]
Saarbrüken 2 14 photon/memory teleportation 2024 [264]
Delft 3 25 memory/memory swapping 2024 [261]
Chattanooga 3 6.7 photon/photon distribution 2024 [302]
Tempe 2 1.6 photon/photon process tomography 2025 [303]